what is the graph of 3x-2y=6
and i need (y) by itself

Found 2 solutions by Fombitz, ewatrrr:
Answer by Fombitz(32388)
Answer by ewatrrr(24785)


Hi,
graphing
3x-2y=6
(3/2)x - 3 = y
the standard slope-intercept form for an equation of a line y = mx + b
where m is the slope and b the y-intercept. slope of this line is (3/2)




Note as y increases/decreases by 3, x increases/decreases by 2